OpenStudy (anonymous):

I will give you a medal if can help me Solve the system using the elimination method. 4/3x+6=-1 4x-4y=13/3 please show work :) Show work PLEASE :)

OpenStudy (anonymous):

i think there is a mistake in the first line

OpenStudy (anonymous):

yes its 6y sorry

OpenStudy (anonymous):

Bribery..

OpenStudy (anonymous):

multiply first equation by 3 to clear the fraction.

OpenStudy (anonymous):

that's my line!

OpenStudy (dumbcow):

x = 3/4 y = -1/3

OpenStudy (anonymous):

thanks can yuo show me how you got it?

OpenStudy (dumbcow):

steps are given in your other posts steps dont change